### v2

# fev-bench: A Realistic Benchmark for Time Series Forecasting

Abstract
Benchmark quality is critical for meaningful evaluation and sustained progress in time series forecasting, particularly with the rise of pretrained models. Existing benchmarks often have limited domain coverage or overlook real-world settings such as tasks with covariates. Their aggregation procedures frequently lack statistical rigor, making it unclear whether observed performance differences reflect true improvements or random variation. Many benchmarks lack consistent evaluation infrastructure or are too rigid for integration into existing pipelines. To address these gaps, we propose fev-bench, a benchmark of 100 forecasting tasks across seven domains, including 46 with covariates. Supporting the benchmark, we introduce fev, a lightweight Python library for forecasting evaluation emphasizing reproducibility and integration with existing workflows. Using fev, fev-bench employs principled aggregation with bootstrapped confidence intervals to report performance along two dimensions: win rates and skill scores. We report results on fev-bench for pretrained, statistical, and baseline models and identify promising future research directions.
Machine Learning, ICML
1 Introduction
Pretrained time series forecasting models are transforming forecasting practice. They often outperform traditional methods (Aksu et al., 2024) while enabling zero-shot inference that simplifies production use and lowers the barrier to entry (Cohen et al., 2025).
Advances in pretrained forecasting models are primarily evaluated on benchmarks, making benchmark quality critical for progress. Shortcomings in existing benchmarks directly shape model development; for example, most general benchmarks ignore covariates despite their prevalence in real-world applications (Bojer & Meldgaard, 2021; Arango et al., 2025). Consequently, most pretrained models lack covariate support, limiting their effectiveness in domains such as retail, where promotional and pricing data are essential for accurate demand forecasting (Fildes et al., 2022).
Beyond task coverage, existing benchmarks often lack statistical rigor. Most studies report single-number summaries, making it unclear whether improvements reflect true advances or random variation. Small gains may vanish or even reverse under minor benchmark changes (Roque et al., 2025), undermining the reliability of conclusions about which models perform better.
Finally, benchmark infrastructure poses additional barriers to progress and reproducibility. Many benchmarks provide only standalone datasets without evaluation code, leading to inconsistent implementations and incomparable results (Hewamalage et al., 2023). When infrastructure exists, it often consists of monolithic systems that bundle models, datasets, and evaluation logic with extensive dependencies, becoming unmaintainable over time. This rigidity prevents extension to new domains or integration into existing workflows, limiting practical utility and longevity.
To address these challenges, we make three contributions:
  • New benchmark. We introduce fev-bench, a forecast evaluation benchmark containing 100 tasks spanning 7 real-world application domains.111huggingface.co/spaces/autogluon/fev-bench Our benchmark addresses a key gap in existing work by including 46 tasks with covariates alongside both univariate and multivariate forecasting scenarios, better reflecting real-world forecasting use cases.
  • Aggregation methods. In our benchmark, we employ principled aggregation strategies including bootstrap-based confidence intervals that quantify whether performance differences are statistically meaningful. This approach enables more reliable model comparisons and assesses the robustness of conclusions to variations in benchmark composition.
  • Evaluation package. We introduce fev, 222github.com/autogluon/fev a lightweight Python package for forecasting evaluation that introduces minimal dependencies while remaining compatible with popular forecasting libraries. The package focuses on reproducibility and extensibility, enabling researchers to easily build and share new benchmarks and the corresponding results.
Benchmark # datasets # tasks # tasks with covariates # multivariate tasks Forecast type
Monash (Godahewa et al., 2021) 42 42 0 0 point
LTSF (Zeng et al., 2023) 9 36 0 9 point
TFB (Qiu et al., 2024) 41 116 0 25 point
BasicTS+ (Shao et al., 2024) 20 40 0 20 point
ProbTS (Zhang et al., 2024) 18 18 0 14 point & quantile
Chronos BM2 (Ansari et al., 2024) 27 27 0 0 point & quantile
GIFT-Eval (Aksu et al., 2024) 55 97 0 43 point & quantile
fev-bench (this work) 96 100 46 35 point & quantile
Table 1: Overview of general time series forecasting benchmarks. The fev benchmark contains more unique datasets than existing benchmarks and includes 46 tasks with covariates, addressing a gap in current evaluation frameworks.
2 Preliminaries
Problem definition. The multivariate time series forecasting problem can be formally stated as follows. We are given a collection $\{{\bm{y}}_{n,1:T}\}_{n=1}^{N}$ of $N$ multivariate time series. For $n=1,\dots,N$ and $t=1,\dots,T$, let ${\bm{y}}_{n,t}=(y_{n,d,t})_{d=1}^{D}\in\mathbb{R}^{D}$ denote the $D$-dimensional observation vector for series $n$ at time $t$, where $D=1$ corresponds to univariate forecasting. The goal is to predict the future $H$ values ${\bm{y}}_{n,T+1:T+H}$ for each series $n$, where $H$ is the forecast horizon. Each time series may be accompanied by covariates $\mathbf{X}_{n,1:T+H}$, including (i) static covariates that do not vary over time (e.g., item or location identifiers), (ii) past-only dynamic covariates observed up to time $T$ (e.g., past related series), and (iii) known dynamic covariates available for all time steps $1,\dots,T+H$ (e.g., holiday indicators).
In the most general form, the goal is to model the conditional distribution $\smash{p\big({\bm{y}}_{n,T+1:T+H}\mid{\bm{y}}_{n,1:T},\mathbf{X}_{n,1:T+H}\big)}$. While full distributional modeling provides the richest information, it is common in practice to produce point forecasts such as conditional means or medians. Alternatively, many applications estimate predictive quantiles $\smash{p(y_{n,d,t}\mid{\bm{y}}_{n,1:T},\mathbf{X}_{n,1:T+H})}$. We denote by ${\mathcal{Q}}\subset(0,1)$ the set of quantile levels and produce forecasts $\smash{\hat{y}^{(q)}_{n,d,t}}$ satisfying $\smash{\Pr(y_{n,d,t}\leq\hat{y}^{(q)}_{n,d,t})=q}$ for all $q\in{\mathcal{Q}}$.
Benchmarks. To evaluate forecasting methods systematically, the general problem must be instantiated as concrete tasks and combined with an evaluation protocol. We refer to such a collection of tasks together with their evaluation and aggregation procedure as a benchmark.
Tasks. A task specifies a concrete forecasting problem. It consists of a dataset and the parameters that define how forecasts are produced and evaluated, including the forecast horizon $H$, evaluation cutoffs, target and covariate columns, and the evaluation metric. A single dataset can yield multiple tasks by varying these parameters. The choice of evaluation metric is integral to task definition, as different metrics correspond to different optimal forecasts; combining conflicting metrics within a single task creates ambiguity about the intended goal (Kolassa, 2020).
Rolling evaluation. Each task is evaluated using a rolling-origin evaluation protocol with $W$ windows (Hyndman & Athanasopoulos, 2018). Let $\tau_{1}<\tau_{2}<\dots<\tau_{W}$ denote the evaluation cutoffs. At each window $w\in\{1,\dots,W\}$, the model receives observations up to $\tau_{w}$ and produces $H$-step forecasts. Advancing the cutoff yields a sequence of forecast–target pairs, resulting in $W$ evaluation windows per task. This setup mimics real-world deployment and provides a more robust estimate of performance over time, especially for datasets with few series.
Aggregation. While tasks define individual evaluation problems, benchmarks must aggregate results across tasks to answer questions such as “Is model A more accurate than model B overall?”. The choice of aggregation directly affects the reliability and interpretability of benchmark results.
We discuss task definitions in Sec. 3, aggregation methods in Sec. 4, and the accompanying software package in Sec. 5.
Benchmark Energy Nature Cloud Mobility Econ Health Retail
GIFT-Eval 16 9 8 7 6 5 4
fev-bench 26 5 20 7 10 8 20
Table 2: Number of datasets from different domains in GIFT-Eval (Aksu et al., 2024) and fev-bench (this work).
Benchmark 10S T 5T 10T 15T 30T H D W M Q Y
GIFT-Eval 2 0 4 2 4 0 13 15 8 5 1 1
fev-bench 0 6 7 2 5 4 22 19 16 7 4 4
Table 3: Number of datasets with different frequencies in GIFT-Eval (Aksu et al., 2024) and fev-bench (this work).
3 Task definitions
We introduce fev-bench (Forecast EValuation Benchmark), a benchmark designed to address limitations of existing forecasting evaluation frameworks. Unlike prior benchmarks that focus on narrow domains or lack evaluation infrastructure, fev-bench offers broad coverage of real-world applications while enabling reproducible and statistically sound comparisons. fev-bench comprises 100 forecasting tasks, with complete specifications in Appendix A. This section describes the design choices guiding its construction.
3.1 Datasets and tasks
Our goal is to construct a general benchmark representative of real-world forecasting applications across domains, frequencies, horizons, and time series characteristics. We consider both univariate and multivariate forecasting problems, with covariates covering dynamic (both past-only and known) and static variables. We evaluate both point and probabilistic forecasting performance.
Datasets. We source time series datasets from established collections including the Monash repository (Godahewa et al., 2021), GIFT-Eval (Aksu et al., 2024), and BOOM (Cohen et al., 2025). However, these collections lack datasets with covariates. To address this gap, we extend the benchmark with public datasets from Kaggle (Bojer & Meldgaard, 2021) and domain-specific repositories (Wang et al., 2023; Lago et al., 2021; Arango et al., 2025).
This curation process yields 96 unique time series datasets.333huggingface.co/datasets/autogluon/fev_datasets We construct 100 forecasting tasks from these datasets using different target column and covariate selections. Among these tasks, 30 include known dynamic covariates, 24 include past dynamic covariates, and 19 include static covariates. These categories are non-exclusive, and both univariate and multivariate tasks may include covariates. Tables 2 and 3 compare our benchmark’s coverage against GIFT-Eval, the most comprehensive existing general forecasting benchmark. fev-bench provides substantially broader coverage across domains and frequencies.
Forecast horizons. Many existing benchmarks reuse the same datasets with different horizons (Zeng et al., 2023; Aksu et al., 2024), creating correlated tasks with limited additional insight. While fev-bench includes both short- and long-horizon tasks, we avoid horizon duplication within datasets. Instead, we select horizons that reflect domain-specific forecasting needs, such as 168 steps for hourly energy demand or 30 steps for daily retail sales.
Rolling evaluation. We use rolling window evaluation to balance computational cost and statistical reliability. The number of windows $W$ depends on dataset size: up to 20 windows for datasets with fewer than 10 series, up to 10 windows for datasets with 10–2000 series, and one window for larger datasets. The value of $W$ is further constrained by data availability, requiring at least $(2\times H+1)$ past observations before the first evaluation window. Results are aggregated across windows using the arithmetic mean.
3.2 Evaluation metrics
Each of the 100 benchmark tasks evaluates point and probabilistic forecast accuracy using complementary metrics. We evaluate point forecast accuracy using Mean Absolute Scaled Error (MASE), following existing benchmarks (Aksu et al., 2024; Godahewa et al., 2021).
$\operatorname{MASE}=\frac{1}{NDH}\sum_{n=1}^{N}\sum_{d=1}^{D}\frac{1}{a_{n,d}}\sum_{t=T+1}^{T+H}|y_{n,d,t}-\hat{y}_{n,d,t}|,$ (1)
where the error for each series $n$ and dimension $d$ is normalized by the historical seasonal error $\smash{a_{n,d}=\frac{1}{T-m}\sum_{t=m+1}^{T}|y_{n,d,t}-y_{n,d,t-m}|}$. Here $\hat{y}_{n,d,t}$ is the point forecast and $m$ is the seasonal period based on the data frequency (e.g., m=12m{=}12 for monthly data). MASE offers several advantages: it is scale-free, balances contributions across series with different magnitudes, handles trends well, and remains robust when the forecast horizon contains zeros (Hyndman & Koehler, 2006).
We use Scaled Quantile Loss (SQL) computed on levels ${\mathcal{Q}}=\{0.1,0.2,...,0.9\}$ for probabilistic forecast accuracy:
$\operatorname{SQL}=\frac{1}{NDH}\sum_{n=1}^{N}\sum_{d=1}^{D}\frac{1}{a_{n,d}}\sum_{t=T+1}^{T+H}\sum_{q\in{\mathcal{Q}}}\rho_{q}(y_{n,d,t},\hat{y}^{(q)}_{n,d,t})$ (2)
where $\hat{y}^{(q)}_{n,d,t}$ is the quantile forecast at level $q$, $a_{n,d}$ is the historical seasonal error as in MASE, and $\rho_{q}(\cdot,\cdot)$ is the quantile loss
[…]
We adopt the Scaled Quantile Loss (SQL) as our primary probabilistic metric, since it is the natural extension of MASE and inherits its desirable scale-independence properties. Still, SQL remains underutilized in forecasting benchmarks, where the scale-dependent Weighted Quantile Loss (WQL) is more common (Ansari et al., 2024; Aksu et al., 2024). Both SQL and WQL are related to the Continuous Ranked Probability Score (CRPS) (Gneiting & Raftery, 2007), Winkler Score and Weighted Interval Score (Tibshirani, 2023), but they differ in how they weight individual series: SQL normalizes by scale, while WQL aggregates in a scale-dependent manner, analogous to the distinction between MASE and Weighted Absolute Percentage Error (WAPE). Our choice follows the same reasoning as the M4 and M5 competitions, which also relied on SQL-equivalent metrics (Makridakis et al., 2020, 2022).
Both MASE and SQL can encounter numerical issues with intermittent time series where the seasonal error $a_{n,d}$ approaches zero (Hewamalage et al., 2023). We verified that this problem does not occur in our benchmark tasks. For completeness, we also report WQL and WAPE scores alongside our primary metrics to enable additional analysis.
3.3 Representative subset of the tasks
In addition to the full fev-bench benchmark consisting of 100 tasks, we provide fev-bench-mini, a curated subset of 20 tasks (Appendix E). These tasks are chosen to capture the diversity of covariates, dimensionalities, domains, and horizons present in the full benchmark, while being small enough to enable rapid iteration and reduced computational cost. fev-bench-mini approximates the relative performance ranking observed on the full benchmark, making it suitable for model development and ablation studies.
4 Aggregating the results
After evaluating $M$ models on $R$ tasks, we obtain the error matrix $E\in\mathbb{R}_{\geq 0}^{R\times M}$, where $E_{rj}$ denotes the error (e.g., MASE) of model $j$ averaged over all evaluation windows of task $r$. Lower values correspond to more accurate forecasts. For models that fail to produce forecasts (e.g., due to timeouts or crashes), we substitute $E_{rj}$ with $E_{r\beta}$, where $\beta$ denotes a predefined baseline model (Seasonal Naive).
4.1 Marginal performance
The primary goal of any benchmark is to rank models by average performance. We use two complementary aggregation methods that capture different aspects of model quality.
Average win rate $W_{j}$ represents the probability that model $j$ achieves lower error than another randomly chosen model $k\neq j$ on a randomly chosen task:
$\begin{split}W_{j}&=\frac{1}{R(M-1)}\sum_{r=1}^{R}\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{M}\Big(\mathds{1}(E_{rj}<E_{rk})\\ &\hskip 85.00012pt+0.5\cdot\mathds{1}(E_{rj}=E_{rk})\Big).\end{split}$ (4)
Here $\mathds{1}(\cdot)$ is the binary indicator function. Ties ($E_{rj}=E_{rk}$) are treated as half-wins for each model. The win rate ranges from 0 (worst) to 1 (best) and provides an intuitive measure of relative model performance. However, win rate has two limitations: it is insensitive to the magnitude of performance differences and changes as new models are added to the benchmark, motivating our second aggregation method.
Skill score (Hyndman & Athanasopoulos, 2018) $S_{j}$ quantifies how much model $j$ reduces forecasting error compared to the fixed baseline model $\beta$ on average:
$S_{j}=1-\sqrt[R]{\prod_{r=1}^{R}\operatorname{clip}\left(\frac{E_{rj}}{E_{r\beta}};\ell,u\right)},$ (5)
where $\operatorname{clip}(x;\ell,u)=\max(\ell,\min(x,u))$ clips $x$ to the interval $\ell,u]$. We aggregate relative errors across tasks using geometric mean, clipping values between $\ell=10^{-2}$ and $u=100$ to avoid excessive influence from extreme values.
Geometric mean aggregation is less sensitive to outliers than the arithmetic mean and ensures that the final ranking remains invariant to the choice of baseline model (Fleming & Wallace, 1986). The geometric mean appropriately handles the multiplicative nature of relative performance comparisons, averaging ratios in a meaningful way where opposing relative errors like $\frac{1}{2}$ and $2$ cancel out.
The skill score ranges from 1 (perfect forecasts) to -$\infty$ (arbitrarily poor performance). Positive values indicate that the model outperforms the baseline on average, while negative values indicate underperformance.
4.2 Pairwise comparison
While marginal performance provides overall rankings, pairwise comparisons reveal specific model relationships that may be obscured in aggregate statistics. The above aggregation methods can be easily generalized to comparing any two models $j$ and $k$.
Pairwise win rate $W_{jk}$ represents the fraction of tasks where model $j$ outperforms model $k$
$W_{jk}=\frac{1}{R}\sum_{r=1}^{R}\left(\mathds{1}(E_{rj}<E_{rk})+0.5\cdot\mathds{1}(E_{rj}=E_{rk})\right).$ (6)
Pairwise skill score $S_{jk}$ quantifies how much model $j$ reduces error compared to model $k$ on average
$S_{jk}=1-\sqrt[R]{\prod_{r=1}^{R}\operatorname{clip}\left(\frac{E_{rj}}{E_{rk}};\ell,u\right)}.$ (7)
4.3 Significance of performance differences
A critical concern in benchmarking is the reliability of reported performance differences. State-of-the-art claims often rest on minor improvements that may vanish under small changes to benchmark composition, casting doubt on whether they reflect genuine advances (Roque et al., 2025).
To address this concern, we compute 95% confidence intervals using paired bootstrap over tasks (Efron, 1992). We generate $B=1000$ bootstrap samples by drawing rows with replacement from $E$, where each $\tilde{E}^{(b)}\in\mathbb{R}_{\geq 0}^{R\times M}$ contains $R$ tasks sampled from the original benchmark. For each $\tilde{E}^{(b)}$, we compute the aggregate statistics to obtain bootstrap distributions such as $\{\tilde{W}_{jk}^{(b)}\}_{b}$. The $(1-\alpha)$ confidence interval for the pairwise win rate $W_{jk}$ is then
$\left[Q_{\nicefrac{{\alpha}}{{2}}}\!\left(\{\tilde{W}_{jk}^{(b)}\}_{b}\right),\;Q_{1-\nicefrac{{\alpha}}{{2}}}\!\left(\{\tilde{W}_{jk}^{(b)}\}_{b}\right)\right],$ (8)
where $Q_{p}(\cdot)$ denotes the empirical $p$-th quantile of the bootstrap distribution. Analogous intervals are constructed for the pairwise skill scores $S_{jk}$. These intervals quantify how conclusions about model comparisons vary under alternative benchmark compositions.
We report bootstrap confidence intervals only for the pairwise statistics ($W_{jk},S_{jk}$), as these directly answer the question of interest: “Does model $j$ consistently outperform model $k$ under different benchmark compositions?” Equivalently, the intervals define paired null-hypothesis tests over tasks, with null hypotheses $W_{jk}=0.5$ (no systematic advantage) and $S_{jk}=0$ (no average error reduction). Confidence intervals for the marginal statistics ($W_{j},S_{j}$) instead describe the variability of each model’s average score in isolation and ignore correlations between models.
Summary. Benchmark interpretation proceeds in two steps. First, the marginal statistics ($W_{j},S_{j}$) provide an overall model ranking. Second, the pairwise statistics with confidence intervals ($W_{jk},S_{jk}$) refine this picture by showing which performance differences are robust to changes in benchmark composition. For example, if a model $j$ ranks highest by marginal win rate $W_{j}$ and all of its pairwise win rates $W_{jk}$ against other models $k\neq j$ have lower bounds above 50%, then model $j$ can be regarded as outperforming every competitor with high confidence.
5 Infrastructure
Comprehensive task coverage and principled evaluation are essential, but standardized infrastructure is equally important for benchmark relevance and longevity. This includes code for task definition, evaluation, and result aggregation.
5.1 Motivation
Existing forecasting benchmarks typically fall into one of two categories: standalone datasets without supporting infrastructure (Godahewa et al., 2021; Zeng et al., 2023), and end-to-end systems that bundle models, datasets, and forecasting tasks (Qiu et al., 2024; Aksu et al., 2024).
Benchmarks consisting only of datasets provide no guarantees that results obtained by different users are comparable. Users may evaluate different forecasting tasks even when referring to the same dataset, due to differences in horizons, cutoffs, metrics, or aggregation strategies.
End-to-end systems address this ambiguity but are often impractical due to their rigidity. They impose many dependencies and assumptions, making extension or integration into existing workflows difficult. Model implementations also tend to become outdated as maintenance costs grow. Version constraints on commonly used libraries such as torch, numpy, and pandas can further reduce transparency. For example, changes in pandas may affect frequency inference, altering the seasonal period $m$ and MASE computation.
5.2 fev Python library
To address these limitations, we introduce fev, a lightweight library that provides essential benchmarking functionality without unnecessary constraints. Its core features include task definition, data loading and splitting, prediction scoring, and result aggregation. fev depends only on Hugging Face datasets (Lhoest et al., 2021) and pydantic (Colvin et al., 2025) for input validation, and does not fix versions of commonly used packages such as torch or numpy. This enables seamless integration into existing pipelines. Using datasets also simplifies data loading and supports extensions such as multimodal forecasting.
fev deliberately excludes model implementations, which are prone to becoming outdated. Instead, it provides adapters that convert data into formats compatible with popular forecasting libraries including GluonTS (Alexandrov et al., 2020), darts (Herzen et al., 2022), StatsForecast (Garza et al., 2022), AutoGluon (Shchur et al., 2023), and sktime (Löning et al., 2019).
Library API. The fev library is built around three main constructs. An EvaluationWindow is a single train–test split defined by a cutoff. A Task specifies a forecasting problem, including the dataset, horizon, covariates, targets, and evaluation metric, and may include multiple evaluation windows. Finally, a Benchmark is a collection of tasks. Benchmarks in fev can be defined using YAML files. Each task produces an evaluation summary that includes both metric values and the full task specification, ensuring unambiguous and comparable results. fev also provides utilities for aggregating results across tasks, as described in Sec. 4.
6 Related work
6.1 Existing benchmarks
Early deep learning approaches for time series forecasting (Lai et al., 2018; Salinas et al., 2020; Rangapuram et al., 2018) did not rely on standardized benchmarks, instead evaluating on a small number of datasets that varied across studies. The M4 (Makridakis et al., 2020) and M5 (Makridakis et al., 2022) competition tasks later gained popularity. The LTSF benchmark (Zhou et al., 2021; Zeng et al., 2023) introduced new datasets with an emphasis on long-horizon forecasting, but remained limited in scope, with a narrow domain focus. It has also been criticized for over-representing similar datasets and using impractical evaluation tasks (Hewamalage et al., 2023).
Subsequent benchmarks, including the Monash repository (Godahewa et al., 2021), BasicTS+ (Shao et al., 2024), TFB (Qiu et al., 2024), ProbTS (Zhang et al., 2024), and Chronos Benchmarks (Ansari et al., 2024), broadened evaluation by covering more domains and supporting both univariate and multivariate forecasting. More recently, GIFT-Eval (Aksu et al., 2024) substantially expanded task diversity across domains and frequencies, becoming a standard benchmark for pretrained time series models. Domain-specific benchmarks such as BOOM (Cohen et al., 2025) have also been introduced. However, none of these benchmarks include forecasting tasks with covariates, despite their practical importance.
In contrast, fev-bench offers broader dataset and domain coverage, including a substantial number of tasks with covariates and domain-appropriate forecasting setups. Complementing this, the fev library provides standardized evaluation infrastructure with principled aggregation and confidence intervals, enabling more robust comparisons.
6.2 Aggregation strategies
Several aggregation methods have been proposed in the literature, each with distinct trade-offs for benchmarking.
Average rank is widely used in forecasting benchmarks (Aksu et al., 2024; Ansari et al., 2024). As shown in Sec. C.1, average rank is mathematically equivalent to the average win rate $W_{j}$, inducing the same model ordering. However, ranks scale with the number of models and do not naturally extend to pairwise comparisons. We therefore prefer win rates, which are bounded between 0 and $1$ and extend directly to pairwise evaluation.
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 91.4 47.3 0.8 0 0
TiRex 82.7 42.6 0.2 1 0
TimesFM-2.5 77.6 42.2 1.9 10 0
Toto-1.0 70.2 40.7 22.1 8 0
Moirai-2.0 64.4 39.3 0.3 28 0
Chronos-Bolt 64.4 38.9 0.2 0 0
TabPFN-TS 63.0 39.6 88.9 0 2
Sundial-Base 46.0 33.4 8.0 1 0
Stat. Ensemble 45.5 20.2 148.6 0 11
Seasonal Naive 20.2 0.0 0.5 0 0
Table 4: Marginal probabilistic forecasting performance of select models (according to the SQL metric) on the full fev-bench benchmark. Results for all models and other metrics are available in Appendix D.
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 88.1 35.5 0.8 0 0
TiRex 76.9 30.0 0.2 1 0
TimesFM-2.5 75.1 30.2 1.9 10 0
Toto-1.0 66.9 28.2 22.1 8 0
Moirai-2.0 61.2 27.3 0.3 28 0
Chronos-Bolt 60.8 26.5 0.2 0 0
TabPFN-TS 58.7 27.6 88.9 0 2
Sundial-Base 53.4 24.7 8.0 1 0
Stat. Ensemble 48.5 15.7 148.6 0 11
Seasonal Naive 21.0 0.0 0.5 0 0
Table 5: Marginal point forecasting performance of select models (according to the MASE metric) on the full fev-bench benchmark. Results for all models and other metrics are available in Appendix D.
Bradley–Terry (Elo) scores (Bradley & Terry, 1952) have been applied across domains, from large language models (Chiang et al., 2024) to tabular benchmarks (Erickson et al., 2025). As shown in Sec. C.2, when all models are compared on all tasks, Bradley–Terry scores induce the same model ranking as average win rates. Thus, average rank, average win rate, and Bradley–Terry scores are equivalent in terms of model ordering under our setup (Sec. 4). We report win rates for simplicity.
Nemenyi post-hoc test with critical difference diagrams (Demšar, 2006) is another rank-based method, as it relies on average ranks. While it controls family-wise error rates, it is often overly conservative and frequently fails to detect meaningful differences (Garcia & Herrera, 2008). Moreover, it provides only binary significance decisions without effect sizes. In contrast, confidence intervals on win rates preserve interpretability while directly conveying the magnitude and uncertainty of performance differences.
Geometric mean relative error (GMRE) (Ansari et al., 2024; Aksu et al., 2024) yields rankings identical to skill scores since $\operatorname{GMRE}_{j}=1-S_{j}$. We adopt this approach with two modifications: clipping extreme values to limit outlier influence, and reporting its complement (the skill score) to maintain a consistent “higher-is-better” interpretation.
7 Results
Figure 1: Pairwise win rates (a) and skill scores (b) of the top-4 models against other models under the SQL metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better. Full pairwise results are available in Appendix D. Best viewed on screen.
In this section we present the experimental evaluation of various forecasting models on the fev-bench benchmark.
7.1 Setup
Models. Our evaluation focuses primarily on pretrained forecasting models that represent the current frontier in time series forecasting research. We select models based on three criteria: strong performance on existing benchmarks such as GIFT-Eval, publicly available implementations, and computational feasibility on consumer hardware (single NVIDIA A10G GPU with 24GB RAM). More details about the model configuration are provided in Appendix B.
We evaluate eight pretrained models: Chronos-2 (Ansari et al., 2025), TimesFM-2.5 (Das et al., 2024), TiRex (Auer et al., 2025), Chronos-Bolt (Base) (Ansari et al., 2024), Toto-1.0 (Cohen et al., 2025), Moirai-2.0 (Woo et al., 2024), TabPFN-TS (Hoo et al., 2025), and Sundial (Liu et al., 2025). Chronos-2 and Toto-1.0 natively support multivariate forecasting ($D>1$); for the remaining models, we convert each multivariate time series into $D$ separate univariate series. Chronos-2 supports both past and future covariates, TabPFN-TS supports known covariates, and the remaining models ignore all covariates.
We include popular statistical baselines: AutoETS, AutoARIMA, and AutoTheta (Garza et al., 2022), as well as the SCUM ensemble combining these univariate models with AutoCES (Petropoulos & Svetunkov, 2020). We also evaluate simple baselines: Seasonal Naive, Naive, and Drift (Hyndman & Athanasopoulos, 2018).
These results represent an initial analysis, and we welcome future submissions from authors of any forecasting approach—pretrained, statistical, or task-specific models.
Evaluation metrics. We follow the evaluation protocol described in Sections 3 and 4. Each task is evaluated using SQL for probabilistic forecasting and MASE for point forecasting, with results aggregated across all 100 tasks using average win rates and skill scores for both marginal and pairwise model comparisons. We also report the following metrics to provide broader context for model performance.
Data leakage. The goal of fev-bench is to assess the zero-shot capabilities of pretrained forecasting models. Two forms of leakage undermine this goal: (i) training on the benchmark training split, which breaks the zero-shot setting, and (ii) exposure to the test split, which constitutes direct test leakage. To prevent both, contributors must indicate, for each model–task pair, whether any part of the dataset at the same frequency was used during training. Resampled variants at different frequencies are not considered leakage. For tasks with reported overlap, we discard the submitted results and impute performance using the strongest fully zero-shot univariate model at release time, Chronos-Bolt (Base). This removes leakage for affected tasks without heavily penalizing the submitted model.
This leakage indicator is required only for pretrained models. Developers may also submit task-specific models, in which case training on the task’s training split is permitted and the zero-shot requirement does not apply. Overall, this policy offers a practical safeguard to separate genuine zero-shot performance from overfitting to benchmark datasets, while recognizing that more subtle forms of leakage may remain difficult to detect.
Runtime. In addition to accuracy indicators, we report the median end-to-end runtime (training plus inference time across all evaluation windows) for each model. Runtimes are normalized to 100 time series per task by scaling each task’s runtime by $100/(N\times W)$, where $N$ is the number of items and $W$ the number of evaluation windows. While this metric is affected by hardware, batch sizes, and implementation details, it still provides useful insights into computational efficiency. In practice, it helps distinguish between modeling paradigms (e.g., sample-based autoregressive models versus direct multi-step forecasters) and between zero-shot and fine-tuned approaches, while incentivizing more efficient implementations.
Model failures. If a model fails to produce a forecast on certain tasks (e.g., due to exceeding the 6 hour runtime limit or encountering an internal error), we replace its performance with the score of the Seasonal Naive baseline.
7.2 Results
Tables 4 & 5 summarize the marginal performance of the top-performing models on fev-bench. Complete results for all models are provided in Appendix D. Chronos-2 achieves the highest average win rate and skill score values, followed by TiRex and TimesFM-2.5. Older pretrained models such as Chronos-Bolt and Sundial rank lower. All pretrained models are substantially more accurate and faster than the best statistical approach (SCUM Ensemble).
To assess whether the observed differences among the leading models reflect genuine improvements rather than evaluation noise, we examine pairwise comparisons under the SQL metric in Figures 1a and 1b. The confidence intervals indicate a statistically significant gap between Chronos-2 and the remaining models. In contrast, TiRex and TimesFM-2.5 show heavy overlap, suggesting no clear winner between the two; under different benchmark compositions or task weightings, either could emerge as the second-best performer.
Figure 2: Average skill scores on the 42 fev-bench tasks with dynamic covariates (based on SQL). Chronos-2 and TabPFN-TS, the only models that use covariates, outperform others on these tasks.
7.3 Directions for improvement
Our evaluation highlights key limitations of current pretrained forecasting models, especially on tasks with covariates and on multivariate forecasting.
Figure 3: Average skill scores relative to the baseline on 35 multivariate tasks of fev-bench (based on SQL). Chronos-2 and Toto-1.0, the only models with native multivariate support, perform well on these tasks.
Forecasting with covariates. On the 42 tasks in fev-bench with dynamic covariates, we compared the same set of pretrained models as in the main evaluation. As shown in Figure 2, models that support covariates (Chronos-2 and TabPFN-TS) achieve higher skill scores on these tasks compared to models that ignore covariates. This demonstrates that covariate support provides meaningful performance gains when such information is available.
Multivariate forecasting. Restricting attention to the 35 multivariate tasks in fev-bench, models with native multivariate support (Chronos-2 and Toto-1.0) perform well (Figure 3). Other models predict each dimension independently. This highlights the benefit of modeling cross-variate dependencies when forecasting multivariate time series, and underscores the need for pretrained models that handle multivariate series directly, though doing so requires advances in both architecture design and access to multivariate training data.
8 Conclusion
We introduced fev-bench, a benchmark that incorporates covariates, multivariate tasks, and principled aggregation methods. By enabling statistically sound comparisons across a broad range of domains, fev-bench provides a reliable foundation for advancing pretrained forecasting models and evaluating their ability to handle real-world requirements such as covariates and multivariate structure.
Complementing this, we presented fev, a lightweight evaluation package designed for reproducibility and extensibility. fev makes it easy to define specialized benchmarks, integrate them into existing workflows, and share results in a consistent way, lowering barriers for future community-driven progress in time series forecasting.
Appendix A Tasks
In total, fev-bench contains 100 time series forecasting tasks. In this section we provide the main statistics of these tasks together with citations to the sources of the datasets. For competition datasets we use their fixed forecast horizon $H$; for all others, $H$ is set by a frequency–horizon mapping, except for a subset of hourly datasets where we use $H=168$ to enable long-range forecasting. The number of evaluation windows $W$ is then chosen to evenly split the series while ensuring that sufficient historical data is available for each forecast of length $H$. Dataset frequencies are reported using pandas frequency aliases (minuTely, Hourly, Daily, Weekly, Monthly, Quarterly, Yearly).
The precise task definitions in YAML format are available under https://github.com/autogluon/fev. The datasets used for evaluation are available under https://huggingface.co/datasets/autogluon/fev_datasets.
A.1 GIFT-Eval
Task Domain Freq. $H$ $W$ Median length # series # targets # past cov. # known cov. # static cov.
BizITObs - L2C cloud 5T 288 20 31,968 1 7 0 0 0
BizITObs - L2C cloud H 24 20 2,664 1 7 0 0 0
ETT energy 15T 96 20 69,680 2 7 0 0 0
ETT energy H 168 20 17,420 2 7 0 0 0
ETT energy D 28 20 724 2 7 0 0 0
ETT energy W 13 5 103 2 7 0 0 0
Hierarchical Sales retail D 28 10 1,825 118 1 0 0 0
Hierarchical Sales retail W 13 10 260 118 1 0 0 0
Hospital healthcare M 12 4 84 767 1 0 0 0
Jena Weather nature 10T 144 20 52,704 1 21 0 0 0
Jena Weather nature D 28 11 366 1 21 0 0 0
Jena Weather nature H 24 20 8,784 1 21 0 0 0
Loop Seattle mobility D 28 10 365 323 1 0 0 0
Loop Seattle mobility 5T 288 10 105,120 323 1 0 0 0
Loop Seattle mobility H 168 10 8,760 323 1 0 0 0
M-DENSE mobility D 28 10 730 30 1 0 0 0
M-DENSE mobility H 168 10 17,520 30 1 0 0 0
SZ Taxi mobility 15T 96 10 2,976 156 1 0 0 0
SZ Taxi mobility H 168 2 744 156 1 0 0 0
Solar energy W 13 1 52 137 1 0 0 0
Solar energy D 28 10 365 137 1 0 0 0
Table 6: Tasks based on datasets coming from the GIFT-Eval corpus (Aksu et al., 2024).
The GIFT-Eval corpus (Aksu et al., 2024) contains various univariate and multivariate datasets, none of which provide covariates. The original datasets have been collected from sources such as Godahewa et al. (2021); Jiang et al. (2023); Mancuso et al. (2021); Wu et al. (2021); Palaskar et al. (2024).
A.2 Macroeconomic datasets
Task Domain Freq. $H$ $W$ Median length # series # targets # past cov. # known cov. # static cov.
Australian Tourism econ Q 8 2 36 89 1 0 0 0
FRED-MD - CEE econ M 12 20 798 1 3 4 0 0
FRED-MD - Macro econ M 12 20 798 1 51 0 0 0
FRED-QD - CEE econ Q 8 20 266 1 3 4 0 0
FRED-QD - Macro econ Q 8 20 266 1 51 0 0 0
GVAR econ Q 8 10 178 33 6 3 0 0
US Consumption econ M 12 10 792 31 1 0 0 0
US Consumption econ Q 8 10 262 31 1 0 0 0
US Consumption econ Y 5 10 64 31 1 0 0 0
World CO2 Emissions econ Y 5 9 60 191 1 0 0 0
World Life Expectancy econ Y 5 10 74 237 1 0 0 0
World Tourism econ Y 5 2 21 178 1 0 0 0
Table 7: Tasks based on various macroeconomic datasets.
We consider various macroeconomic datasets such as GVAR (Mohaddes & Raissi, 2024), US Consumption (Wilms & Croux, 2016), Australian Tourism (Athanasopoulos et al., 2009), FRED-MD (McCracken & Ng, 2016), FRED-QD (McCracken & Ng, 2020), world CO2 emmissions (Kag, 2025a), life expectancy (Kag, 2025b) and global tourism (Kag, 2025d).
For each of FRED-MD and FRED-QD, we create two forecasting tasks. The first follows the CEE model (Christiano et al., 1999) and focuses on forecasting employment, inflation, and federal funds rate indicators. The second task involves jointly forecasting 51 core macroeconomic indicators. Note that we use the snapshot of FRED-MD corresponding to August 2025, which is different from FRED-MD snapshot used in Godahewa et al. (2021).
A.3 Energy datasets
Task Domain Freq. $H$ $W$ Median length # series # targets # past cov. # known cov. # static cov.
ENTSO-e Load energy 15T 96 20 175,292 6 1 0 3 0
ENTSO-e Load energy 30T 96 20 87,645 6 1 0 3 0
ENTSO-e Load energy H 168 20 43,822 6 1 0 3 0
EPF-BE energy H 24 20 52,416 1 1 0 2 0
EPF-DE energy H 24 20 52,416 1 1 0 2 0
EPF-FR energy H 24 20 52,416 1 1 0 2 0
EPF-NP energy H 24 20 52,416 1 1 0 2 0
EPF-PJM energy H 24 20 52,416 1 1 0 2 0
ERCOT energy D 28 20 6,452 8 1 0 0 0
ERCOT energy H 168 20 154,872 8 1 0 0 0
ERCOT energy M 12 15 211 8 1 0 0 0
ERCOT energy W 13 20 921 8 1 0 0 0
GFC12 energy H 168 10 39,414 11 1 0 1 0
GFC14 energy H 168 20 17,520 1 1 0 1 0
GFC17 energy H 168 20 17,544 8 1 0 1 0
Solar with Weather energy 15T 96 20 198,600 1 1 2 7 0
Solar with Weather energy H 24 20 49,648 1 1 2 7 0
Table 8: Tasks based on datasets related to energy generation and consumption.
These datasets include electricity price forecasting (EPF) benchmark (Lago et al., 2021), ERCOT generation data (Ansari et al., 2024), ENTSO-e load data (Open Power System Data, 2020) with weather originating from Renewables.ninja (Staffell et al., 2023), and solar generation with weather (Kag, 2025c).
A.4 BOOMLET
Task Domain Freq. $H$ $W$ Median length # series # targets # past cov. # known cov. # static cov.
BOOMLET - 1062 cloud 5T 288 20 16,384 1 21 0 0 0
BOOMLET - 1209 cloud 5T 288 20 16,384 1 53 0 0 0
BOOMLET - 1225 cloud T 60 20 16,384 1 49 0 0 0
BOOMLET - 1230 cloud 5T 288 20 16,384 1 23 0 0 0
BOOMLET - 1282 cloud T 60 20 16,384 1 35 0 0 0
BOOMLET - 1487 cloud 5T 288 20 16,384 1 54 0 0 0
BOOMLET - 1631 cloud 30T 96 20 10,463 1 40 0 0 0
BOOMLET - 1676 cloud 30T 96 20 10,463 1 100 0 0 0
BOOMLET - 1855 cloud H 24 20 5,231 1 52 0 0 0
BOOMLET - 1975 cloud H 24 20 5,231 1 75 0 0 0
BOOMLET - 2187 cloud H 24 20 5,231 1 100 0 0 0
BOOMLET - 285 cloud T 60 20 16,384 1 75 0 0 0
BOOMLET - 619 cloud T 60 20 16,384 1 52 0 0 0
BOOMLET - 772 cloud T 60 20 16,384 1 67 0 0 0
BOOMLET - 963 cloud T 60 20 16,384 1 28 0 0 0
Table 9: Tasks based on datasets from BOOMLET (Cohen et al., 2025).
We include the multivariate observability datasets from the BOOMLET benchmark (Cohen et al., 2025). BOOMLET is a subset of the larger BOOM benchmark curated by the original authors. We additionally limit our attention to datasets with frequency of at least 1 minute to avoid including too many datasets from a single source to fev-bench.
A.5 Forecasting competitions
Task Domain Freq. $H$ $W$ Median length # series # targets # past cov. # known cov. # static cov.
Favorita Store Sales retail M 12 2 54 1,579 1 1 1 6
Favorita Store Sales retail W 13 10 240 1,579 1 1 1 6
Favorita Store Sales retail D 28 10 1,688 1,579 1 1 2 6
Favorita Transactions retail M 12 2 54 51 1 1 0 5
Favorita Transactions retail W 13 10 240 51 1 1 0 5
Favorita Transactions retail D 28 10 1,688 51 1 1 1 5
KDD Cup 2022 energy D 14 10 243 134 1 9 0 0
KDD Cup 2022 energy 10T 288 10 35,279 134 1 9 0 0
KDD Cup 2022 energy 30T 96 10 11,758 134 1 9 0 0
M5 retail M 12 1 58 30,490 1 0 8 5
M5 retail W 13 1 257 30,490 1 0 8 5
M5 retail D 28 1 1,810 30,490 1 0 8 5
Restaurant retail D 28 8 296 817 1 0 0 4
Rohlik Orders retail W 8 5 170 7 1 9 4 0
Rohlik Orders retail D 61 5 1,197 7 1 9 4 0
Rohlik Sales retail W 8 1 150 5,243 1 1 13 7
Rohlik Sales retail D 14 1 1,046 5,390 1 1 13 7
Rossmann retail W 13 8 133 1,115 1 1 4 10
Rossmann retail D 48 10 942 1,115 1 1 5 10
Walmart retail W 39 1 143 2,936 1 0 10 4
Table 10: Tasks based on datasets coming from various forecasting competitions
We use datasets from forecasting competitions held on kaggle.com (Bojer & Meldgaard, 2021). These include Favorita store sales & transactions (Kag, 2020), the M5 competition (Makridakis et al., 2022), restaurant visitor & reservation (Kag, 2017), Rossmann (Kag, 2015), Walmart (Kag, 2014), and Rohlik (Roh, 2025) store sales forecasting competitions. We also consider the KDD Cup 2022 dataset where the goal is to predict wind power generation (Zhou et al., 2022), and the Global Energy Forecasting Competitions held in 2012, 2014 and 2017 (Hong et al., 2014).
A.6 Other sources
Task Domain Freq. $H$ $W$ Median length # series # targets # past cov. # known cov. # static cov.
ECDC ILI healthcare W 13 10 201 25 1 0 0 0
Hermes retail W 52 1 261 10,000 1 0 1 2
Hospital Admissions healthcare D 28 20 1,731 8 1 0 0 0
Hospital Admissions healthcare W 13 16 246 8 1 0 0 0
Redset cloud 5T 288 10 25,920 118 1 0 0 1
Redset cloud 15T 96 10 8,640 126 1 0 0 1
Redset cloud H 24 10 2,160 138 1 0 0 1
UCI Air Quality nature H 168 20 9,357 1 4 0 3 0
UCI Air Quality nature D 28 11 389 1 4 0 3 0
UK COVID - Nation - Cumulative healthcare D 28 20 729 4 3 5 0 0
UK COVID - Nation - Cumulative healthcare W 8 4 105 4 3 5 0 0
UK COVID - Nation - New healthcare D 28 20 729 4 3 5 0 0
UK COVID - Nation - New healthcare W 8 4 105 4 3 5 0 0
UK COVID - UTLA - Cumulative healthcare W 13 5 104 214 1 0 0 0
UK COVID - UTLA - New healthcare D 28 10 721 214 1 0 0 0
Table 11: Tasks based on datasets collected from other sources.
We also include datasets from the following miscellaneous sources.
  • Influenza-like-illness cases collected by the European Centre for Disease Prevention and Control (ECDC, 2025).
  • Fashion trend data from Hermes (David et al., 2022).
  • Hospital admissions data from Riyadh (of Health Affairs & Ministry of Health, 2025).
  • Query counts for Amazon Redshift database servers (van Renen et al., 2024).
  • Solar energy generation with corresponding weather covariates (Kag, 2025c).
  • Air quality measurements in an Italian city with accompanying weather data (Vito et al., 2008).
  • COVID-19 cases, hospital admissions, and deaths in the United Kingdom at different administrative levels (Kag, 2025e).
Appendix B Models
We evaluate seven pretrained models on fev-bench, whose key properties are summarized in Table 12. Most are decoder-only transformers, except TiRex, a decoder-only xLSTM model, Chronos-Bolt (Base), an encoder-decoder transformer, and Chronos-2, an encoder-only model. All models except TabPFN-TS process non-overlapping patches of time series rather than individual observations. Toto, TabPFN-TS, and Sundial produce sample forecasts, while the others generate quantiles on a fixed grid. Toto and Chronos-2 are the only models that natively supports multivariate targets. TabPFN-TS supports known covariates, though we did not provide past covariates since they consistently degraded its accuracy. Chronos-2 supports both past and known covariates.
For all pretrained models, we keep hyperparameters at their default values unless specified in the table. For Toto, we reduced the samples per batch and batch size to avoid out-of-memory errors on large multivariate datasets. We run all pretrained models on a g5.2xlarge AWS instance with a single A10G GPU (24GB GPU RAM, 32GB RAM), using PyTorch 2.6 with CUDA 12.6 via AWS Deep Learning Containers. The version in the table below refers to either the PyPI package version, or the date on which the official repository was cloned if no PyPI package is provided by the authors.
Model Name Hugging Face ID Batch size Version Max Context Hyperparameters
Chronos-2 amazon/chronos-2 100 2.2.2 8192 {cross_learning: True}
TiRex NX-AI/TiRex 512 2025-09-01 2048 -
Toto Datadog/Toto-Open-Base-1.0 24 2025-08-01 4096 {samples_per_batch: 8}
Moirai 2.0 Salesforce/moirai-2.0-R-small 128 2025-08-10 4000 -
Chronos-Bolt amazon/chronos-bolt-base 256 1.5.3 2048 -
TimesFM 2.5 google/timesfm-2.5-200m-pytorch 256 2025-09-28 16000 -
TabPFN-TS Prior-Labs/TabPFN-v2-reg - 1.0.3 5000 {checkpoint: ’2noar4o2’}
Sundial thuml/sundial-base-128m 512 2025-09-01 2880 -
Table 12: Properties of different pretrained time series forecasting models.
We also include statistical baselines from the StatsForecast library (Garza et al., 2022), such as AutoETS, AutoARIMA, AutoTheta, as well as the SCUM ensemble (Petropoulos & Svetunkov, 2020). To avoid long runtimes, we truncate the context length of statistical models to 1000 steps and set the maximum season length to 200 for AutoETS, AutoTheta, AutoARIMA and SCUM Ensemble. For evaluation of statistical models we used StatsForecast v2.0.1 and ran the experiments on the m6i.4xlarge AWS instances with 16 vCPU cores and 64GB RAM.
Appendix C Extended discussion of aggregation methods
This appendix clarifies how average win rates $W_{j}$ (Eq. 4) relate to other aggregation methods commonly used in benchmarking, such as average ranks (Aksu et al., 2024; Ansari et al., 2024) and Bradley–Terry (“Elo”) scores (Erickson et al., 2025). We show that all three induce the same ordering of models.
C.1 Average win rate and average rank are equivalent
For a given task $r$ and model $j$, let
$M_{\text{lower}}=\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{M}\mathds{1}(E_{rk}<E_{rj}),\qquad M_{\text{tied}}=\sum_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{M}\mathds{1}(E_{rk}=E_{rj}).$
The midrank of $j$ on task $r$ is
$\operatorname{rank}_{rj}\;=\;1+M_{\text{lower}}+\tfrac{1}{2}M_{\text{tied}}.$
Its contribution to $W_{j}$ equals
$\frac{1}{M-1}\sum_{k\neq j}\Big(\mathds{1}(E_{rj}<E_{rk})+\tfrac{1}{2}\,\mathds{1}(E_{rj}=E_{rk})\Big)=1-\frac{\operatorname{rank}_{rj}-1}{M-1}.$
Averaging over tasks gives
$W_{j}=1-\frac{\overline{\operatorname{rank}}_{j}-1}{M-1},\qquad\overline{\operatorname{rank}}_{j}=\tfrac{1}{R}\sum_{r=1}^{R}\operatorname{rank}_{rj}.$
Thus, $W_{j}$ and $\overline{\operatorname{rank}}_{j}$ are affinely equivalent and induce the same ordering of models (with lower rank $\leftrightarrow$ higher win rate).
C.2 Average win rate and Bradley–Terry (Elo) scores result in the same ranking
The Bradley–Terry (BT) model, also known as Elo rating (Chiang et al., 2024), provides a parametric way to convert pairwise win rates into latent skill scores. In contrast to the nonparametric average win rate $W_{j}$, the BT model assumes that each model $j$ has an underlying skill parameter $\theta_{j}\in\mathbb{R}$, and that the probability of $j$ outperforming $k$ follows a logistic link:
$\Pr(E_{rj}<E_{rk})\;=\;\sigma\!\big(\lambda(\theta_{j}-\theta_{k})\big),\qquad\sigma(x)=\tfrac{1}{1+e^{-x}},\;\;\lambda>0.$
Here $\lambda$ is a scaling constant (in Elo, $\lambda=\ln 10/400$). The parameters $\theta=(\theta_{1},\dots,\theta_{M})$ are estimated by maximum likelihood:
$\hat{\theta}\;\in\;\arg\max_{\theta\in\mathbb{R}^{M}}\sum_{j<m}\Big[W_{jm}\log\sigma(\lambda(\theta_{j}-\theta_{m}))+(1-W_{jm})\log\sigma(\lambda(\theta_{m}-\theta_{j}))\Big].$
Typically some identifiability constraint is added, such as fixing $\theta_{\beta}=1000$ for a chosen baseline $\beta$.
Proposition C.1.
Suppose all $M$ models are compared on the same $R$ tasks, with pairwise win rates $W_{jk}$ (Eq. 6) and average win rates $W_{j}=\tfrac{1}{M-1}\sum_{k\neq j}W_{jk}$ (Eq. 4). At the BT MLE with scale $\lambda>0$,
$\theta_{j}>\theta_{k}\;\;\Longleftrightarrow\;\;W_{j}>W_{k},\qquad\theta_{j}=\theta_{k}\;\;\Longleftrightarrow\;\;W_{j}=W_{k}.$
Proof.
Differentiating the log-likelihood gives the score equations
$\frac{\partial\ell}{\partial\theta_{j}}=\lambda\sum_{m\neq j}\Big(W_{jm}-\sigma(\lambda(\theta_{j}-\theta_{m}))\Big)=0.$
Subtracting the equations for $j$ and $k$ yields
$(M-1)(W_{j}-W_{k})=\sum_{m\neq j,k}\big[\sigma(\lambda(\theta_{j}-\theta_{m}))-\sigma(\lambda(\theta_{k}-\theta_{m}))\big]+\big[\sigma(\lambda\Delta)-\sigma(-\lambda\Delta)\big],$
where $\Delta=\theta_{j}-\theta_{k}$. Each term on the right is strictly increasing in $\Delta$, so the whole expression has the same sign as $\Delta$. Thus $\operatorname{sign}(W_{j}-W_{k})=\operatorname{sign}(\theta_{j}-\theta_{k})$, with equality iff $\Delta=0$. Strict concavity of the BT log-likelihood ensures uniqueness of the solution up to translation. ∎
Conclusion. Average win rate $W_{j}$ and BT/Elo scores $\theta_{j}$ induce the same ordering of models, with higher win rate corresponding to higher Elo score.
Appendix D Extended results
Tabular results for each dataset-task combination, updated marginal and pairwise results in an interactive format are available under https://huggingface.co/spaces/autogluon/fev-bench.
The failures for Stat. Ensemble, AutoARIMA, and AutoETS models in the tables below correspond to the models exceeding the 6 hour time limit for a single task. TabPFN-TS failed on 2 tasks due to out of memory errors.
D.1 Marginal performance
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 91.4 47.3 0.8 0 0
TiRex 82.7 42.6 0.2 1 0
TimesFM-2.5 77.6 42.2 1.9 10 0
Toto-1.0 70.2 40.7 22.1 8 0
Moirai-2.0 64.4 39.3 0.3 28 0
Chronos-Bolt 64.4 38.9 0.2 0 0
TabPFN-TS 63.0 39.6 88.9 0 2
Sundial-Base 46.0 33.4 8.0 1 0
Stat. Ensemble 45.5 20.2 148.6 0 11
AutoARIMA 40.7 20.6 19.5 0 10
AutoETS 33.7 -26.8 3.5 0 3
AutoTheta 27.1 5.5 3.3 0 0
Seasonal Naive 20.2 0.0 0.5 0 0
Naive 13.8 -45.4 0.5 0 0
Drift 9.3 -45.8 0.5 0 0
Table 13: Marginal probabilistic forecasting performance of all models (according to the SQL metric) on the full fev-bench benchmark. The reported metrics are defined in Sections 4.1 and 7.1.
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 88.1 35.5 0.8 0 0
TiRex 76.9 30.0 0.2 1 0
TimesFM-2.5 75.1 30.2 1.9 10 0
Toto-1.0 66.9 28.2 22.1 8 0
Moirai-2.0 61.2 27.3 0.3 28 0
Chronos-Bolt 60.8 26.5 0.2 0 0
TabPFN-TS 58.7 27.6 88.9 0 2
Sundial-Base 53.4 24.7 8.0 1 0
Stat. Ensemble 48.5 15.7 148.6 0 11
AutoARIMA 36.7 11.2 19.5 0 10
AutoTheta 34.9 11.0 3.3 0 0
AutoETS 33.2 2.3 3.5 0 3
Seasonal Naive 21.0 0.0 0.5 0 0
Naive 19.3 -16.7 0.5 0 0
Drift 15.4 -18.1 0.5 0 0
Table 14: Marginal point forecasting performance of all models (according to the MASE metric) on the full fev-bench benchmark. The reported metrics are defined in Sections 4.1 and 7.1.
D.2 Pairwise comparison
Figure 4: Pairwise win rates $W_{jk}$ (Eq. 6) of all models against each other under the scaled quantile loss (SQL) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better.
Figure 5: Pairwise skill scores $S_{jk}$ (Eq. 7) of all models against each other under the scaled quantile loss (SQL) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better. Note that pairwise skill score is not symmetric, $S_{jk}\neq S_{kj}$.
Figure 6: Pairwise win rates $W_{jk}$ (Eq. 6) of all models against each other under the mean absolute scaled error (MASE) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better.
Figure 7: Pairwise skill scores $S_{jk}$ (Eq. 7) of all models against each other under the mean absolute scaled error (MASE) metric on fev-bench, with 95% confidence intervals obtained via bootstrapping. Higher values are better. Note that pairwise skill score is not symmetric, $S_{jk}\neq S_{kj}$.
Appendix E fev-bench-mini: A representative subset of fev-bench
E.1 Tasks
fev-bench-mini consists of 20 tasks that are representative of the full 100 tasks comprising fev-bench (Appendix A).
Task Domain Freq. $H$ $W$ Median length # series # targets # past cov. # known cov. # static cov.
BOOMLET - 1282 cloud T 60 20 16,384 1 35 0 0 0
BOOMLET - 1676 cloud 30T 96 20 10,463 1 100 0 0 0
BOOMLET - 619 cloud T 60 20 16,384 1 52 0 0 0
BizITObs - L2C cloud 5T 288 20 31,968 1 7 0 0 0
EPF-NP energy H 24 20 52,416 1 1 0 2 0
ETT energy 15T 96 20 69,680 2 7 0 0 0
ETT energy H 168 20 17,420 2 7 0 0 0
GFC14 energy H 168 20 17,520 1 1 0 1 0
Hospital Admissions healthcare W 13 16 246 8 1 0 0 0
Hospital Admissions healthcare D 28 20 1,731 8 1 0 0 0
Jena Weather nature H 24 20 8,784 1 21 0 0 0
M-DENSE mobility D 28 10 730 30 1 0 0 0
Rohlik Orders retail D 61 5 1,197 7 1 9 4 0
Rossmann retail W 13 8 133 1,115 1 1 4 10
Rossmann retail D 48 10 942 1,115 1 1 5 10
Solar with Weather energy H 24 20 49,648 1 1 2 7 0
UCI Air Quality nature H 168 20 9,357 1 4 0 3 0
UK COVID - Nation - Cumulative healthcare D 28 20 729 4 3 5 0 0
US Consumption econ Y 5 10 64 31 1 0 0 0
World CO2 Emissions econ Y 5 9 60 191 1 0 0 0
Table 15: Tasks included in fev-bench-mini.
E.2 Evaluation results
For completeness, we provide the evaluation results on fev-bench-mini. The overall ranking of the models, win rates and skill scores align with the scores on the full benchmark reported in Appendix D.
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 91.8 51.1 1.2 0 0
TiRex 80.4 43.4 0.4 0 0
TimesFM-2.5 74.8 44.0 5.2 5 0
Toto-1.0 74.1 43.0 44.3 5 0
TabPFN-TS 67.1 46.2 176.3 0 0
Moirai-2.0 65.0 41.7 0.4 30 0
Chronos-Bolt 61.4 40.4 0.5 0 0
Sundial-Base 50.0 37.6 8.3 0 0
Stat. Ensemble 45.7 27.3 261.8 0 2
AutoARIMA 44.3 30.6 79.6 0 2
AutoETS 32.9 -9.4 6.4 0 0
AutoTheta 25.7 8.5 5.1 0 0
Seasonal Naive 16.4 0.0 0.8 0 0
Naive 12.9 -36.9 0.8 0 0
Drift 7.5 -36.0 0.8 0 0
Table 16: Marginal probabilistic forecasting performance of all models (according to the SQL metric) on the fev-bench-mini benchmark. The reported metrics are defined in Sections 4.1 and 7.1.
Model Avg. win rate (%) Skill score (%) Median runtime / 100 series (s) Leakage (%) # failures
Chronos-2 88.9 40.4 1.2 0 0
TiRex 75.0 31.4 0.4 0 0
TimesFM-2.5 73.7 32.3 5.2 5 0
Toto-1.0 72.3 30.9 44.3 5 0
TabPFN-TS 63.6 35.1 176.3 0 0
Moirai-2.0 62.1 30.1 0.4 30 0
Chronos-Bolt 59.6 28.4 0.5 0 0
Sundial-Base 58.6 29.8 8.3 0 0
Stat. Ensemble 48.2 21.1 261.8 0 2
AutoARIMA 41.4 20.4 79.6 0 2
AutoTheta 33.2 14.8 5.1 0 0
AutoETS 31.8 2.3 6.4 0 0
Seasonal Naive 15.0 0.0 0.8 0 0
Naive 13.9 -17.4 0.8 0 0
Drift 12.5 -17.2 0.8 0 0
Table 17: Marginal point forecasting performance of all models (according to the MASE metric) on the fev-bench-mini benchmark. The reported metrics are defined in Sections 4.1 and 7.1.